# Language Model Distillation: A Temporal Difference Imitation Learning Perspective

### v1

@@ 5 Experiments
5.1 Experiment Setup
We take instruction-following [44] as the conditional text generation task, where models are trained to generate responses according to the instructions. We fine-tune a large model on a dataset consisting of instruction-response pairs as the teacher model. Then, we compare different KD methods by evaluating the student model’s instruction-following performance.
Model selections. We conduct experiments using three model families: the GPT-2 family [45], the OPT family [46], and the Qwen-2.5 family [6]. For the GPT-2 family, we use the 1.5B model as the teacher and the 120M, 340M, and 760M models as students. In the OPT family, the 6.7B model serves as the teacher, with the 125M, 350M, and 1.3B models as students. For the Qwen-2.5 family, we use the 3B model as the teacher and the 0.5B model as the student.
Training. We build our training dataset using databricks-dolly-15K (https://github.com/databrickslabs/dolly/tree/master), which contains 15,000 human-authored instruction-response pairs. To accommodate model constraints, we remove any samples that exceed the maximum context length. From the remaining data, we randomly select 500 examples for validation and 1,000 for testing, leaving around 12,500 samples for training. For the pretraining dataset $\mathcal{D}^{PT}$, we adopt OpenWebText [47] for models in the GPT-2 and Qwen-2.5 families, and use the RoBERTa training corpus [48] for OPT models. Following the MiniLLM setup [19], we select hyperparameters based on Rouge-L [49] scores evaluated on the validation set.
@@ 5.1 Experiment Setup
Baselines choices. Following [19], we choose four baselines: (i) SFT directly fine-tunes the student model on $\mathcal{D}$ with golden responses; (ii) KD [9, 50] fine-tunes the student model on $\mathcal{D}$ using the teacher distribution as supervision at each token step, also known as word-level KD; (iii) SeqKD [8, 51, 52, 53, 54] fine-tunes the student model on the response sequences generated by the teacher model; (iv) MiniLLM [19] fine-tunes the student model using a policy gradient approach, where the reward is defined by the reverse KL divergence between the output distributions of the teacher and student.
Evaluation. Following [19], we evaluate the trained models on three instruction-following datasets: (i) DollyEval: The 500-sample test set we split from the databricks-dolly-15K dataset; (ii) SelfInst: [55] A user-oriented instruction-following set with 252 samples; (iii) Vicuna: [51] The 80 challenging questions used in the Vicuna evaluation.
Metrics. Following [19], we employ two complementary evaluation metrics: (i) Rouge-L [49] to assess the precision of generated responses, following prior work demonstrating its effectiveness for large-scale instruction-following evaluation [56]; (ii) Win rate of our approach evaluated by a GPT-4o-mini [57] oracle against baseline methods to measure generation quality.
@@
Figure 2: Comparison of win rates against KD, SeqKD, and MiniLLM baselines. We evaluate using GPT-4o-mini [57] as the judging oracle, with Qwen-2.5 (3B) as the teacher model and a smaller (0.5B) model as the student. Results are based on 500 responses per distilled model, generated under the Dolly evaluation setting.
5.2 Experimental Results
[…]
@@ 5.2 Experimental Results
5.3 Impact of $p$
5.4 Training Time
Online vs offline training. In general, the choice between online and offline training is a trade-off between computational cost and performance. While auto-regressive online generation is computationally expensive, offline training allows for the reuse of pre-collected datasets, making it more efficient. However, online training typically yields better final performance [58, 59, 60, 38]. In particular, Wen et al. [20], Ko et al. [21] discussed how the offline setting can accelerate distillation, especially given the high cost of online generation for LLMs. We choose offline training due to computational constraints, although the proposed approach is inherently agnostic to the choice of training setting.
Wall-time comparison. To demonstrate the computational efficiency of our offline approach, we compare its training time to reach the optimal validation performance with the online method Gu et al. [19]. All experiments are conducted on 4×4$\times$4 ×A40 GPUs. Results are shown in Table 3.
5.5 Ablation on the $\chi^{2}$ Regularization
We perform an ablation study on the use of $\chi^{2}$ regularization, as introduced in Section , to assess its impact on distillation performance. Our empirical results show that incorporating $\chi^{2}$ regularization improves the performance of the distilled student model. The study is conducted on two distillation setups: GPT-2 1.5B to 760M and Qwen-2.5 3B to 0.5B. The results are presented in Table 4.
Table 3: Training time comparison: Our approach achieves the optimal validation performance in significantly less training time compared to Gu et al. [19], thanks to its offline training paradigm.
@@ 5.5 Ablation on the $\chi^{2}$ Regularization
6 Discussions
Further related works. From the imitation learning (IL) perspective, language model distillation can be viewed as a form of IL—or more broadly, reinforcement learning (RL)—in large discrete action spaces [61, 62, 63, 64, 65], a significant line of research in deep (inverse) RL. A common approach in this setting is to develop methods that restrict the action space to a manageable subset. This remains a core challenge in (inverse) RL, where little prior knowledge is available about the full action space. In contrast, distillation benefits from access to a teacher model, which enables efficient identification of high-quality action subsets. This insight directly motivates our top-$p$ TD learning design.
Limitations. A limitation of this work—and of white-box language model distillation in general—is the requirement that the teacher and student share an identical vocabulary. This constraint arises from the need to match their output distributions, either explicitly (as in BC) or implicitly (as in IRL).
Conclusions. To summarize, motivated by the observation that a small number of top tokens contribute the majority of the total probability mass, we proposed a top-$p$ TD learning framework that operates over a top-$p$ candidate action subset $\mathcal{A}_{p}^{\star}$. This framework is, in principle, plug-and-play and can be integrated into any TD-based IRL algorithm. We demonstrated its practicality through a concrete implementation of top-$p$ IQL and showed its empirical performance gains.
@@ A.2 Proposition 2
For brevity, we drop the norm subscripts $\infty$ and $\mathcal{A}_{p}^{\star}$, and we use $\hat{Q}_{p}^{\star}$ and ^$\hat{\mathcal{B}}^{\star}_{p}$ to denote $\bar{Q}^{\textsc{proj}_{p}\pi^{\star}}$ and $\mathcal{B}^{\textsc{proj}_{p}\pi^{\star}}_{p}$, respectively.
@@
since ^$\hat{\mathcal{B}}^{\star}_{p}$ is a contractor,
@@ Appendix B Training Details
Across all settings, our BD distillation process uses a batch size of 64 and a learning rate of 5e-6. We set $p=0.8$ for all experiments, except for the OPT 6.7B to 1.3B distillation, where we use $p=0.5$. We use the regularization strength $\alpha=0.1$ and the discount factor $\gamma=0.99$ for all settings. Following the setup of Gu et al. [19], we also adopt a two-phase training strategy for our approach:
@@
  • Phase 2: We then further fine-tune the initialized student model using our distillation algorithm 1 on the teacher-generated dataset $\mathcal{D}^{\star}$. The student model is trained for 3 epochs, and we select the checkpoint with the best validation Rouge-L score. All experiments are conducted on a machine with 4×4$\times$4 × A40 GPUs.
@@ Appendix C Data Generation and Evaluation Details
During data generation and evaluation, we sample from each model using a temperature of 1, conditioned on a given query from the corresponding dataset. For evaluation, we use five random seeds—$[10,20,30,40,50]$—to perform sampling. During data generation, we sample 8 responses from the teacher model for each query in the prompt dataset to construct the teacher offline dataset $\mathcal{D}_{T}$. To convert instruction-response pairs into complete sentences, we apply the prompt wrapper shown in Figure 3.
@@
For the win rate evaluation, we use the prompt wrapper shown in Figure 4 to format the instruction along with one response from our method and one from a baseline method into a single sentence, which is then evaluated by the GPT-4o-mini [57] oracle. To mitigate potential bias from the ordering of options, we conduct each comparison twice—once with our method as option A and the baseline as option B, and once with the roles reversed. We then average the win rates from both evaluations.
@@
Figure 4: Prompt wrapper for win rate evaluation using GPT-4o-mini [57].

### v2

@@ Introduction
Experiment Setup
We take instruction-following (ouyang2022training) as the conditional text generation task, where models are trained to generate responses according to the instructions. We fine-tune a large model on a dataset consisting of instruction-response pairs as the teacher model. Then, we compare different KD methods by evaluating the student model’s instruction-following performance.
Model selections. We conduct experiments using three model families: the GPT-2 family (radford2019language), the OPT family (zhang2022opt), and the Qwen-2.5 family (yang2024qwen2). For the GPT-2 family, we use the 1.5B model as the teacher and the 120M, 340M, and 760M models as students. In the OPT family, the 6.7B model serves as the teacher, with the 125M, 350M, and 1.3B models as students. For the Qwen-2.5 family, we use the 3B model as the teacher and the 0.5B model as the student.
Training. We build our training dataset using databricks-dolly-15K (https://github.com/databrickslabs/dolly/tree/master), which contains 15,000 human-authored instruction-response pairs. To accommodate model constraints, we remove any samples that exceed the maximum context length. From the remaining data, we randomly select 500 examples for validation and 1,000 for testing, leaving around 12,500 samples for training. For the pretraining dataset $\mathcal{D}^{PT}$, we adopt OpenWebText (Gokaslan2019OpenWeb) for models in the GPT-2 and Qwen-2.5 families, and use the RoBERTa training corpus (liu2019roberta) for OPT models. Following the MiniLLM setup (gu2023minillm), we select hyperparameters based on Rouge-L (lin-2004-rouge) scores evaluated on the validation set.
@@
Baselines choices. Following (gu2023minillm), we choose four baselines: (i) SFT directly fine-tunes the student model on $\mathcal{D}$ with golden responses; (ii) KD (sanh2019distilbert; song2020lightpaff) fine-tunes the student model on $\mathcal{D}$ using the teacher distribution as supervision at each token step, also known as word-level KD; (iii) SeqKD (kim2016sequence; vicuna2023; alpaca; peng2023instruction; zhou2023lima) fine-tunes the student model on the response sequences generated by the teacher model; (iv) MiniLLM (gu2023minillm) fine-tunes the student model using a policy gradient approach, where the reward is defined by the reverse KL divergence between the output distributions of the teacher and student.
Evaluation. Following (gu2023minillm), we evaluate the trained models on three instruction-following datasets: (i) DollyEval: The 500-sample test set we split from the databricks-dolly-15K dataset; (ii) SelfInst: (wang2022self) A user-oriented instruction-following set with 252 samples; (iii) Vicuna: (vicuna2023) The 80 challenging questions used in the Vicuna evaluation.
Metrics. Following (gu2023minillm), we employ two complementary evaluation metrics: (i) Rouge-L (lin-2004-rouge) to assess the precision of generated responses, following prior work demonstrating its effectiveness for large-scale instruction-following evaluation (wang2022super); (ii) Win rate of our approach evaluated by a GPT-4o-mini (openai2024gpt4omini) oracle against baseline methods to measure generation quality.
@@
Table 2: Impact of $p$: We demonstrate the impact of different choices of $p$ across three models from distinct model families. Results are reported using the Rouge-L score and averaged over five random seeds. The performance gains relative to the baseline (p= 1.0p\,{=}\,1.0, without top-$p$ mask) are shown in green. It is evident that our top-$p$ framework achieves consistent improvements.
@@
Table 3: Training time comparison: Our approach achieves the optimal validation performance in significantly less training time compared to gu2023minillm, thanks to its offline training paradigm.
@@
Experimental Results
Impact of $p$
Training Time
Online vs offline training. In general, the choice between online and offline training is a trade-off between computational cost and performance. While auto-regressive online generation is computationally expensive, offline training allows for the reuse of pre-collected datasets, making it more efficient. However, online training typically yields better final performance (klein2011batch; lee2019truly; jarrett2020strictly; garg2021iq). In particular, wen2023f; ko2024distillm discussed how the offline setting can accelerate distillation, especially given the high cost of online generation for LLMs. We choose offline training due to computational constraints, although the proposed approach is inherently agnostic to the choice of training setting.
Wall-time comparison. To demonstrate the computational efficiency of our offline approach, we compare its training time to reach the optimal validation performance with the online method gu2023minillm. All experiments are conducted on $4\timesA$40 GPUs. Results are shown in Table 3.
Ablation on the $\chi^{2}$ Regularization
We perform an ablation study on the use of $\chi^{2}$ regularization, as introduced in the implementation section, to assess its impact on distillation performance. Our empirical results show that incorporating $\chi^{2}$ regularization improves the performance of the distilled student model. The study is conducted on two distillation setups: GPT-2 1.5B to 760M and Qwen-2.5 3B to 0.5B. The results are presented in Table 4.
Conclusion
Our work connects language model distillation to imitation learning in large discrete action spaces, where prior methods struggle without action space priors, while distillation benefits from teacher guidance—motivating our top-$p$ TD learning approach. A key limitation is the shared vocabulary requirement between teacher and student for distribution matching. We propose a plug-and-play top-$p$ TD framework that focuses on high-probability tokens, demonstrating empirical gains when integrated with IQL.
Technical Appendix
@@ Appendix A Omitted Proofs
For brevity, we drop the norm subscripts $\infty$ and $\mathcal{A}_{p}^{\star}$, and we use $\hat{Q}_{p}^{\star}$ and $\hat{\mathcal{B}}^{\star}_{p}$ to denote $\bar{Q}^{\textsc{proj}_{p}\pi^{\star}}$ and $\mathcal{B}^{\textsc{proj}_{p}\pi^{\star}}_{p}$, respectively.
@@
since $\hat{\mathcal{B}}^{\star}_{p}$ is a contractor,
@@ Appendix B Training Details
Across all settings, our BD distillation process uses a batch size of 64 and a learning rate of 5e-6. We set $p=0.8$ for all experiments, except for the OPT 6.7B to 1.3B distillation, where we use $p=0.5$. We use the regularization strength $\alpha=0.1$ and the discount factor $\gamma=0.99$ for all settings. Following the setup of gu2023minillm, we also adopt a two-phase training strategy for our approach:
@@
  • Phase 2: We then further fine-tune the initialized student model using our distillation algorithm 1 on the teacher-generated dataset $\mathcal{D}^{\star}$. The student model is trained for 3 epochs, and we select the checkpoint with the best validation Rouge-L score. All experiments are conducted on a machine with $4\times$ A40 GPUs.
@@ Appendix C Data Generation and Evaluation Details
During data generation and evaluation, we sample from each model using a temperature of 1, conditioned on a given query from the corresponding dataset. For evaluation, we use five random seeds—[10,20,30,40,50][10,20,30,40,50]—to perform sampling. During data generation, we sample 8 responses from the teacher model for each query in the prompt dataset to construct the teacher offline dataset $\mathcal{D}_{T}$. To convert instruction-response pairs into complete sentences, we apply the prompt wrapper shown in Figure 3.
@@
For the win rate evaluation, we use the prompt wrapper shown in Figure 4 to format the instruction along with one response from our method and one from a baseline method into a single sentence, which is then evaluated by the GPT-4o-mini (openai2024gpt4omini) oracle. To mitigate potential bias from the ordering of options, we conduct each comparison twice—once with our method as option A and the baseline as option B, and once with the roles reversed. We then average the win rates from both evaluations.
@@
Figure 4: Prompt wrapper for win rate evaluation using GPT-4o-mini (openai2024gpt4omini).